# FinResearchBench: A Logic Tree based Agent-as-a-Judge Evaluation Framework for Financial Research Agents

### v2

@@ Abstract.
Recently, AI agents are rapidly evolving in intelligence and widely used in professional research applications, such as STEM, software development, finance, etc. Among these AI agents, deep research agent is a key category as it can perform long-horizon tasks and solve problems of greater complexity. However, there are few evaluation frameworks and benchmarks that systematically and automatically investigate the capabilities of these research agents. Furthermore, financial research problems have distinct complexity and subtlety. To fill in the gap, we propose FinResearchBench, which is a logic tree based Agent-as-a-Judge and targets specifically for the financial research agents. It provides a comprehensive and automatic assessment of the research agents across 7 key types of tasks in the financial research domain. The contributions of this work are two-folded: (1) the first and innovative Agent-as-a-Judge system that extracts the logic tree of the research outcome and uses it as the intermediate information to present a comprehensive, reliable and robust evaluation; (2) finance oriented that it covers 70 typical financial research questions, spreading across 7 frequently-encountered types of tasks in the domain.
@@
copyright: acmlicensedjournalyear: 2018doi: XXXXXXX.XXXXXXXconference: Make sure to enter the correct conference title from your rights confirmation email; June 03–05, 2018; Woodstock, NYisbn: 978-1-4503-XXXX-X/2018/06ccs: Computing methodologies Natural language generationccs: Applied computing Economics
@@ 1. Introduction
Large Language Models (LLMs) have been rapidly developing in recent years and demonstrate superior performance in diverse tasks, such as story creation (Zhao et al., 2023), machine translation (Xu et al., 2024), dialogue system (Yi et al., 2024), sentiment analysis (Zhang et al., 2023b), etc. Furthermore, LLMs are dramatically evolving in both long text understanding and generation (Team et al., 2024; Anthropic, 2024). The intrinsic capability of LLM to generate long text is often limited, which is revealed to be bounded by the training data the model has seen during the pre-training and alignment stages (Bai et al., 2024). Yet, long text generation is essential, especially in professional research areas, such as literature, law, STEM and finance. Numerous methods have been proposed to improve LLM’s competence in long text generation, including scaling the generation length by adding long output training data (Bai et al., 2024), performing generation in a multi-agent framework (Hong et al., 2023), setting up as deep research agents (Schmidgall et al., 2025; Zheng et al., 2025), and so on. Deep research agent is a prominent category of AI agents and is getting more and more important, since it could generate well-articulated, source-attributed research report by autonomously browsing the internet, calling various tools, collecting information from different sources, etc. The research result is of high value and brings huge productivity improvement. Therefore, comprehensive evaluation of these deep research agents becomes critical. Various benchmarks and evaluation methods are proposed, such as HelloBench (Que et al., 2024), Deep Research Bench (Bosse et al., 2025), LongEval (Wu et al., 2025), etc. However, these methods do not fully utilize the results of the intermediate steps, which could be the deciding factor for the quality of the research outcome. Furthermore, existing benchmarks and evaluation methods are of general purpose and face severe challenges when applied for financial research, since it requires deep understanding of finance-specific terminologies, clauses and domain-specific insights. In this work, we propose FinResearchBench, a logic tree based Agent-as-a-Judge that is designed specifically for financial research agents. It first extracts the logic tree from the research outcome and then uses the extracted logic tree to build an agentic system to evaluate the financial research agents. Moreover, FinResearchBench is finance concentrated and covers 7 key types of tasks in the financial research domain. With the aim of the extracted logic tree as the intermediate step, FinResearchBench is able to provide a comprehensive, reliable and robust evaluation for various financial research agents.
@@ 2.1. LLM and Deep Research Agent
Large Language Models (LLMs) have shown remarkable performance in various tasks, such as story creation, machine translation, dialogue system, etc (Zhao et al., 2023; Xu et al., 2024; Yi et al., 2024; Zhang et al., 2023b). In terms of dealing with long text, they also demonstrate superior capability in understanding and processing (Team et al., 2024; Anthropic, 2024). However, the capability of generating long text progresses rather slowly. Many AI agent systems are then developed including single-agent and multi-agent ones. And it represents a paradigm shift from standalone LLMs towards more autonomous, task-oriented systems (Hong et al., 2023; Wu et al., 2023). AI agents largely increase the upper bound of the generation results since they are able to perform more actions, like web browsing, tool calls, code execution, etc. Additionally, AI agents are capable of processing the obtained information more effectively and accurately, such as Chain-of-Throughts (CoT) method (Wei et al., 2022), Tree-of-Thoughts (ToT) method (Yao et al., 2023a), ReAct method (Yao et al., 2023b), etc. Deep research agent is a growing category of AI agents, since it can orchestrate multiple steps in an autonomous manner, including internet browsing, information retrieval, tool calling, etc. It is able to provide in-depth research of highly complicated problems, and thus brings huge productivity improvement and high commercial values.
@@ 2.2. General and Domain-Specific Evaluation of Deep Research Agent
The evaluation of AI agents is critical since it determines if the emergent agents are applicable for real-world problems or not. Moreover, the evaluation will guide the direction for further improvement. Numerous general benchmarks and evaluation methods are proposed, such as AgentBench (Liu et al., 2023), ToolQA (Zhuang et al., 2023), HelloBench (Que et al., 2024), DeepResearch Bench (Du et al., 2025), LongEval (Wu et al., 2025), etc. These benchmarks are constructed with diverse tools in different environments, enabling extensive evaluation of given agents. For deep research agent, it is a subcategory of the AI agents and performs multi-step orchestration in an autonomous manner. It produces research report that contains rich information and is well articulated. Many evaluation methods are proposed in the literature, including Agent-as-a-Judge (Zhuge et al., 2024), Mind2Web (Deng et al., 2023), etc.
However, these methods are majorly targeted for general domains and inadequate for vertical ones, especially knowledge-heavy domains like law, medicine, STEM and finance. In order to close the gap, many domain-specific benchmarks and evaluation methods are introduced, for instance LegalAgentBench (Li et al., 2024a), literature review evaluation (Tang et al., 2024), FinEval (Zhang et al., 2023a) and xbench (Chen et al., 2025). These domain-specific benchmarks and evaluations are designed to offer valuable and domain-specific insights for corresponding research.
@@ 3.1.1. Logic Tree
In order to draft a comprehensive and insightful professional research document in finance or any other domain, it is mandatory to do a thorough literature review, prepare ideas, and collect supporting evidence before starting to write the full report. Cognitive and linguistic theory defines different stages, i.e. Pre-Writing, Writing and Re-Writing. It defines the planning stage that contains all the preparation steps (Rohman, 1965). It lays the foundation for a potentially coherent, insightful generation outcome. In this paper, we define the underlying structure of ideas, arguments, evidences, etc. as the logic tree.
It is natural to derive that discovering the underlying logic tree of any financial research outcome and using it as the intermediate step will reduce the difficulty of the evaluation. Furthermore, the logic tree has some quantifiable metrics, such as depth, width, number of nodes, etc. Consequently, rule-based metrics can be formulated, such as analysis width, analysis depth, information density, as listed in Table 1.
@@ 3.1.4. Logic Tree Based Agent-as-a-Judge
Leveraging the extracted logic tree as the intermediate step, FinResearchBench is developed. It is an agentic system and essentially a logic tree based Agent-as-a-Judge, as illustrated in Figure 2. With the rich intermediate information throughout the writing process, FinResearchBench is able to provide comprehensive, reliable and robust evaluation in an automatic way. As indicated in Table 1, there are several metrics that are rule derived, either based on the extracted logic tree or not. We explain the rule-based metrics as follows. Notations used in the following section are detailed in Appendix B.
@@ 3.2. Data Construction
Financial research is skewed and the highly-encountered issues are closely related to monetary returns, such as financial plate investigation, stock-specific analysis, etc. Therefore, there is a huge potential for productivity improvement and real-world values for an accurate, comprehensive and reliable evaluation of these agents. Additionally, financial research agents have some unique intricacies, such as domain-specific terms, clauses and in-depth knowledge. Therefore, it is crucial to collect sufficiently complex, yet practically applicable financial research tasks.
  • Query Sampling: Randomly sample user queries from our online product: AI Cashcat (https://www.cashcat.cn). A strict privacy protection process is enforced to all raw queries before subsequent steps, including anonymization, personal identifiable information removal, session information disposal, etc.
@@
  • Data Curation: Conduct interviews with domain experts from finance news agencies, research department of security companies, professional investors, etc., and generate 70 questions, 10 typical research problems per category.
@@ 4.1.2. Agents and Experiments
We conduct a two-stage evaluation. In the first stage, we run the FinResearchBench on a range of leading international and domestic AI agents as well as our proprietary product, Cashcat Pro ( https://cashcat.cn/?mode=DeepResearch). Cashcat Pro is a deep research agent developed by FinStep, targeted specifically for the finance sector. It integrates a multi-agent framework for collaborative analysis, a suite of professional tools for real-time financial data, and cutting-edge AI models to deliver deep, actionable insights and automate the entire research process.
@@ 4.2. Main Results
    Not officially accessible in Chinese Mainland.
@@
    Not officially accessible in Chinese Mainland.
@@ 4.3.6. Insights and Conclusions
We perform a correlation analysis between the total number of nodes from the extracted logic tree and the GT. As detailed in the above table, the analysis yields a Spearman correlation coefficient of $r=0.48$, with a highly statistically significant p-value of less than 0.00010.00010.0001.

### v3

@@ Abstract.
Recently, AI agents are rapidly evolving in intelligence and are widely used in professional research applications, such as STEM, software development, and finance. Among these AI agents, deep research agent is a key category as it can perform long-horizon tasks and solve problems of greater complexity. However, there are few evaluation frameworks and benchmarks that systematically and automatically investigate the capabilities of these research agents. In addition, financial research problems have distinct complexity and subtlety. To fill in the gap, we propose FinResearchBench, which is a logic tree-based Agent-as-a-Judge and targets specifically for the financial research agents. It provides a comprehensive and automatic assessment of the research agents across 7 key types of tasks in the financial research domain. The contributions of this work are two-folded: (1) the first and innovative Agent-as-a-Judge system that extracts the logic tree of the research outcome and uses it as the intermediate information to present a comprehensive, reliable, and robust evaluation; (2) finance-oriented that it covers 70 typical financial research questions, spreading across 7 frequently-encountered types of task in the domain.
@@
copyright: acmlicensedjournalyear: 2025copyright: acmlicensedconference: 6th ACM International Conference on AI in Finance; November 15–18, 2025; Singapore, Singaporebooktitle: 6th ACM International Conference on AI in Finance (ICAIF ’25), November 15–18, 2025, Singapore, Singaporedoi: 10.1145/3768292.3770364isbn: 979-8-4007-2220-2/2025/11
@@ 1. Introduction
Large Language Models (LLMs) have been rapidly developing in recent years and demonstrate superior performance in diverse tasks, such as story creation (Zhao et al., 2023), machine translation (Xu et al., 2024), dialogue system (Yi et al., 2025), and sentiment analysis (Zhang et al., 2024). Furthermore, LLMs are dramatically evolving in both long text understanding and generation (Team et al., 2024; Anthropic, 2024). The intrinsic capability of LLM to generate long text is often limited, which is revealed to be bounded by the training data the model has seen during the pretraining and alignment stages (Bai et al., 2024). Yet, long text generation is essential, especially in professional research areas, such as literature, law, STEM, and finance. Numerous methods have been proposed to improve LLM’s competence in long text generation, including scaling the generation length by adding long output training data (Bai et al., 2024), performing generation in a multi-agent framework (Hong et al., 2024), and setting up as deep research agents (Schmidgall et al., 2025; Zheng et al., 2025). Deep research agent is a prominent category of AI agents and is getting more and more important, since it could generate a well-articulated, source-attributed research report by autonomously browsing the Internet, calling various tools, collecting information from different sources. The research result is highly valuable and brings a huge improvement in productivity. Therefore, a comprehensive evaluation of these deep research agents becomes critical. Various benchmarks and evaluation methods are proposed, such as HelloBench (Que et al., 2024), Deep Research Bench (FutureSearch et al., 2025), and LongEval (Wu et al., 2025). However, these methods do not fully utilize the results of the intermediate steps, which could be the deciding factor for the quality of the research outcome. Furthermore, existing benchmarks and evaluation methods are of general purpose and face severe challenges when applied for financial research, since it requires deep understanding of finance-specific terminologies, clauses and domain-specific insights. In this work, we propose FinResearchBench, a logic tree based Agent-as-a-Judge that is designed specifically for financial research agents. It first extracts the logic tree from the research outcome and then uses the extracted logic tree to build an agentic system to evaluate the financial research agents. Moreover, FinResearchBench is finance concentrated and covers 7 key types of tasks in the financial research domain. With the aim of the extracted logic tree as the intermediate step, FinResearchBench is able to provide a comprehensive, reliable and robust evaluation for various financial research agents.
@@ 2.1. LLM and Deep Research Agent
Large Language Models (LLMs) have shown remarkable performance in various tasks, such as story creation, machine translation, and dialogue system (Zhao et al., 2023; Xu et al., 2024; Yi et al., 2025; Zhang et al., 2024). In terms of dealing with long text, they also demonstrate superior capability in understanding and processing (Team et al., 2024; Anthropic, 2024). However, the capability of generating long text progresses rather slowly. Many AI agent systems are then developed including single-agent and multi-agent ones. And it represents a paradigm shift from standalone LLMs towards more autonomous, task-oriented systems (Hong et al., 2024; Wu et al., 2023). AI agents largely increase the upper bound of the generation results since they are able to perform more actions, like web browsing, tool calls, and code execution. Additionally, AI agents are capable of processing the obtained information more effectively and accurately, such as Chain-of-Throughts (CoT) method (Wei et al., 2022), Tree-of-Thoughts (ToT) method (Yao et al., 2023a), and ReAct method (Yao et al., 2023b). Deep research agent is a growing category of AI agents, since it can orchestrate multiple steps in an autonomous manner, including internet browsing, information retrieval, and tool calling. It is able to provide in-depth research of highly complicated problems, and thus brings huge productivity improvement and high commercial values.
@@ 2.2. General and Domain-Specific Evaluation of Deep Research Agent
The evaluation of AI agents is critical since it determines if the emergent agents are applicable for real-world problems or not. Moreover, the evaluation will guide the direction for further improvement. Numerous general benchmarks and evaluation methods are proposed, such as AgentBench (Liu et al., 2024), ToolQA (Zhuang et al., 2023), HelloBench (Que et al., 2024), DeepResearch Bench (Du et al., 2025), and LongEval (Wu et al., 2025). These benchmarks are constructed with diverse tools in different environments, enabling extensive evaluation of given agents. For deep research agent, it is a subcategory of the AI agents and performs multi-step orchestration in an autonomous manner. It produces research report that contains rich information and is well articulated. Many evaluation methods are proposed in the literature, including Agent-as-a-Judge (Zhuge et al., 2024), and Mind2Web (Deng et al., 2023).
However, these methods are majorly targeted for general domains and inadequate for vertical ones, especially knowledge-heavy domains like law, medicine, STEM and finance. In order to close the gap, many domain-specific benchmarks and evaluation methods are introduced, for instance LegalAgentBench (Li et al., 2024a), literature review evaluation (Tang et al., 2024), FinEval (Guo et al., 2024) and xbench (Chen et al., 2025). These domain-specific benchmarks and evaluations are designed to offer valuable and domain-specific insights for corresponding research.
@@ 3.1.1. Logic Tree
In order to draft a comprehensive and insightful professional research document in finance or any other domain, it is mandatory to do a thorough literature review, prepare ideas, and collect supporting evidence before starting to write the full report. Cognitive and linguistic theory defines different stages, i.e. Pre-Writing, Writing and Re-Writing. It defines the planning stage that contains all the preparation steps (Rohman, 1965). It lays the foundation for a potentially coherent, insightful generation outcome. In this paper, we define the underlying structure of ideas, arguments, and evidences as the logic tree.
It is natural to derive that discovering the underlying logic tree of any financial research outcome and using it as the intermediate step will reduce the difficulty of the evaluation. Furthermore, the logic tree has some quantifiable metrics, such as depth, width, and number of nodes. Consequently, rule-based metrics can be formulated, such as analysis width, analysis depth, information density, as listed in Table 1.
@@ 3.1.4. Logic Tree Based Agent-as-a-Judge
Leveraging the extracted logic tree as the intermediate step, FinResearchBench is developed. It is an agentic system and essentially a logic tree based Agent-as-a-Judge, as illustrated in Figure 2. With the rich intermediate information throughout the writing process, FinResearchBench is able to provide comprehensive, reliable and robust evaluation in an automatic way. As indicated in Table 1, there are several metrics that are rule derived, either based on the extracted logic tree or not. We explain the rule-based metrics as follows. Notations used in the following section are detailed in Appendix B. We note that the hyper-parameters in these scoring functions (e.g., constants and weights) have been determined empirically. This approach is taken to strike a balance between the various analytical dimensions, such as width and depth, while ensuring the resulting scores are sufficiently discriminative.
@@ 3.2. Data Construction
Financial research is skewed and the highly-encountered issues are closely related to monetary returns, such as financial plate investigation and stock-specific analysis. Therefore, there is a huge potential for productivity improvement and real-world values for an accurate, comprehensive and reliable evaluation of these agents. Additionally, financial research agents have some unique intricacies, such as domain-specific terms, clauses and in-depth knowledge. Therefore, it is crucial to collect sufficiently complex, yet practically applicable financial research tasks.
  • Query Sampling: Randomly sample user queries from our online product: AI Cashcat (https://www.cashcat.cn). A strict privacy protection process is enforced to all raw queries before subsequent steps, including anonymization, personal identifiable information removal and session information disposal.
@@
  • Data Curation: Conduct interviews with domain experts from finance news agencies, research department of security companies, professional investors, and generate 70 questions, 10 typical research problems per category.
@@ 4.1.2. Agents and Experiments
We conduct a two-stage evaluation. In the first stage, we run the FinResearchBench on a range of leading international and domestic AI agents as well as our proprietary product, Cashcat Pro ( https://cashcat.cn/?mode=DeepResearch). Cashcat Pro is a deep research agent developed by FinStep, targeted specifically for the finance sector. It integrates a multi-agent framework for collaborative analysis, a suite of professional tools for real-time financial data, and cutting-edge AI models to deliver deep, actionable insights and automate the entire research process. In addition, Gemini is powered by Gemini Pro 2.5, while OpenAI DeepResearch is powered by GPT-4o.
@@ 4.2. Main Results
*Not officially accessible in Chinese Mainland.
@@
*Not officially accessible in Chinese Mainland.
@@ 4.3.6. Insights and Conclusions
We perform a correlation analysis between the total number of nodes from the extracted logic tree and the GT. As detailed in the above table, the analysis yields a Spearman correlation coefficient of $r=0.48$, with a highly statistically significant p-value of less than $0.0001$.
@@ 5.2. Challenges and Future Works
Moving from Logic Trees to Logic Graphs: Furthermore, a more fundamental limitation lies in the choice of a tree structure itself to represent a document’s logic. While trees excel at capturing hierarchical, top-down arguments, they inherently struggle to model the complex, network-like nature of sophisticated reasoning. Real-world argumentation is rarely a strict hierarchy; it often involves cross-references where a single piece of evidence supports multiple disparate claims, feedback loops, or premises that jointly support a conclusion in a non-linear fashion.
A logic graph, particularly a directed acyclic graph (DAG), offers a far more robust and complete representational framework. Such a structure can naturally capture these many-to-many relationships, allowing for a more nuanced and accurate evaluation of a document’s overall coherence and logical integrity. Therefore, a pivotal direction for future work is to move beyond tree-based models and develop novel methods for extracting and analyzing these more comprehensive logic graphs. This would represent a significant step towards an AI that can truly understand the intricate web of arguments presented in a document.
Acknowledgements.
This paper was supported by ”2025 Special Program for Supporting Innovative Development in Leading Industries (AI Track) under the High-Quality Industrial Development Initiative” (Project Name: Finstep Finsmart Intelligent Service Platform; Project ID: 2025-GZL-RGZN-01024) from Shanghai Municipal Commission of Economy and Informatization, Shanghai, China.
@@ Appendix B Notations
Appendix C Example Prompt
This appendix provides illustrative examples of the prompt used to evaluate the reports.
C.1. Logic Tree Extraction
The prompt used to extract the logic tree.
Listing 1: Prompt for Logic Tree Extraction
1"""
2##BACKGROUND
3Youareanexpertinlogicalanalysis,specializinginidentifyingandconstructinglogicaltreesfromtext.Toevaluatethequalityofgivenanswersbasedonmetricssuchasinformationvolume,logicalcoherence,depth,andbreadth,youmustfirstidentifythecorearguments,sub-arguments,andevidencewithinthetextandconstructalogicaltreetorepresentthem.
4##TASK
5Youwillreceiveaqueryandthecorrespondinggeneratedanswer.Thisanswercouldbeacasualreply,aresearchsummaryforaconcept,oraprofessionalresearchreport.Yourtaskistocarefullyreadtheanswer,identifyallcorearguments,sub-arguments,andevidence,andconstructalogicaltreetorepresenttheirrelationships.
6Thecoreargumentservesastherootnodeofthelogicaltree,sub-argumentsareintermediatenodes,andevidenceformstheleafnodes.Allroot,intermediate,andleafnodesmustbeextracteddirectlyfromtheprovidedtext.Youmustnotaddcontentnotmentionedintheoriginaltextormakesubjectiveinferences.
7##DEFINITIONS
81.␣␣**Evidence**:Objectivefactspresentedinthetext,includingbutnotlimitedtostatementsofobjectivesituations,events,ordata.
92.␣␣**Sub-argument**:Aviewpointoropinionderivedfromevidenceorothersub-argumentsthroughrigorouslogicalreasoning.
103.␣␣**CoreArgument**:Acentralclaimderivedfromevidenceand/orsub-argumentsthatcondensestheessenceofasignificantportion,ortheentirety,ofthetext.Itrepresentsthemaintakeawayofthereport.
11##PRINCIPLESFORCONSTRUCTINGTHELOGICALTREE
121.␣␣**IdentifyEvidence**:Readthetextcarefullytoidentifyeverypieceofevidence.Asinglesentencemaycontainmultiplepiecesofevidence,oronepieceofevidencemaybeonlyapartofasentence.Striveforexhaustiveidentificationwithoutomission.
132.␣␣**ConstructEvidenceNodes**:Whencreatingevidencenodes,youmust**strictlyadheretotheinformationprovidedintheoriginaltext**.Donotaddextraneousinformation,makesubjectiveinferences,oraltertheoriginalmeaning.
143.␣␣**ConstructIntermediateNodes(Sub-arguments)**:
15␣␣␣␣*Asub-argumentcanbeformedinoneofthreeways:
16␣␣␣␣␣␣␣␣(i)Derivedentirelyfromevidence(leafnodes),representingaconclusiondrawndirectlyfromfacts.Thesearetypicallyfoundclosetotheleavesofthetree.
17␣␣␣␣␣␣␣␣(ii)Derivedfromacombinationofevidenceandothersub-arguments.Thesearecommoninthemiddlelayersofthetree.
18␣␣␣␣␣␣␣␣(iii)Derivedentirelyfromothersub-arguments,representingahigher-levelconclusion.Thesearetypicallyfoundclosertotheroot.
19␣␣␣␣*Alldeductionsmustbelogicallysound,withaclearandexplicitrelationshipbetweenaparentnodeanditschildren.
204.␣␣**ConstructtheRootNode(CoreArgument)**:Therootnodeisprimarilyderivedfromsub-arguments,sometimessupplementedbykeyevidence.
21
22UserQuery:
23{query}
24ArticleContent:
25{report}
26
27##OUTPUTFORMAT
28PleasestrictlyadheretothefollowingJSONformat.Note:
291.AllstringsmustbeenclosedinEnglishdoublequotes("").
302.Thereshouldbenoquotationmarkswithinastringvalue.
313.EnsuretheJSONsyntaxiscorrect,includingcommas,brackets,andbraces.
324.Donotincludeanycommentsoradditionalexplanationsintheoutput.
335.Thetreedoesnotneedtobestrictlythreelevelsdeep;itsdepthandbreadtharedeterminedbythearticle’sstructure.
346.Ifthearticlecontainsmultipleindependentcorearguments,therootnodecanbenull,andthecoreargumentscanbetreatedasfirst-levelchildren.
357.Generally,thecomplexityofthelogicaltreeshouldbeproportionaltothelengthandcomplexityoftheanswer.Keepthisinmindduringextraction.
36##REQUIREMENTS
371.␣␣Identifythecoreargument(s)ofthetext(rootnode).
382.␣␣Identifythesub-argumentsandevidencethatsupportthecoreargument(s)(childnodes).
393.␣␣Statementssupportedbydataareclassifiedasevidence;allothersareconsideredarguments.
404.␣␣Leafnodescanbeeitherargumentsorevidence,dependingonthespecificcontextandtheruleabove.
41"""